# Hitting probability for Reflected Brownian Motion at Small Target

### v2

@@ 6. Proof of Main Result in General SULD Domains
As we remarked we consider the case when $\epsilon$ is very small. In this case we can localize our domain
@@ 6.1. Localization of Domain
Observe that when the RBM particle is near the target the hitting probability is asymptotically irrelevant to the shape of domain.
Without loss of generality let $\mathrm{tg}\in U^{1}$ and let $\mathscr{C}_{1}$, $\mathscr{C}_{2}$ be two fixed $n-1$ dimensional subsets in $U^{1}$ such that $\Phi_{1}(\mathscr{C}_{1})=\partial B_{P}(r_{1})\cap\Phi_{1}(U_{1})$ and $\Phi_{1}(\mathscr{C}_{2})=\partial B_{P}(r_{2})\cap\Phi_{1}(U_{1})$ for some $r_{1}>r_{2}>0$ and that both curves do not intersects $\mathrm{abs}$. As both $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$ are disjoint compact sets their distance is separated and we may assume that any curve $\gamma:[0,1]\to\Omega$ connecting some points on $\mathrm{abs}$ and $P=\gamma(1)$ that only intersects $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$ once for each, intersects $\mathscr{C}_{1}$ at $t_{1}$ and $\mathscr{C}_{2}$ at $t_{2}$ where $t_{1}<t_{2}$ ($\mathscr{C}_{2}$ lies “closer” to $P$ than $\mathscr{C}_{1}$). We let $\tau_{1}=\inf\{t\geq 0:B_{t}^{x_{0}}\in\mathscr{C}_{2}\}$ for $i=1,2$. Let $\mathbb{P}_{1}:=\mathbb{P}(B^{x_{0}}_{\tau}\in\mathrm{tg},B^{x_{0}}_{[\tau_{1},\tau]}\cap\mathscr{C}_{1}=\emptyset)$. We claim the following observation:
@@
There exists constants $c_{1}$, $c_{2}$ such that
@@
Since our RBM particle roams outside $\Phi_{1}^{-1}(\overline{B_{P}(r_{1})}\cap\Phi_{1}(U_{1}))$ we have a smaller chance to hit $\mathrm{tg}$ so $\mathbb{P}_{1}\geq\mathbb{P}(B^{x_{0}}_{\tau}\in\mathrm{tg},B^{x_{0}}_{[\tau_{1},\tau]}\cap\mathscr{C}_{1}\neq\emptyset)$. Thus $\mathbb{P}\left(B^{x_{0}}_{\tau}\in\mathrm{tg}\right)\leq 2\mathbb{P}_{1}$. ∎
@@
(3) $\mathbb{P}\left(B^{x_{0}}_{\tau}\in\mathrm{tg}\right)=\mathbb{P}(\tau_{1}\leq\infty)\times\mathbb{P}\left(B_{\tau}^{B^{x_{0}}_{\tau_{1}}}\in\mathrm{tg}\right)$
as any RBM hits target will hits $\mathscr{C}_{1}$ by geometrical observation. Now from above proposition for asymptotic behavior we may consider $\mathbb{P}\left(B_{\tau}^{B^{x_{0}}_{\tau_{1}}}\in\mathrm{tg},B^{x_{0}}_{[\tau_{1},\tau]}\cap\mathscr{C}_{1}=\emptyset\right)$. This probability can be interpreted as the probability that the RBM starts at a point $\mathscr{C}_{2}$, roaming in $\Phi_{1}^{-1}(\overline{B_{P}(r_{1})}\cap\Phi_{1}(U_{1}))$ and hits target, which is same as the target hitting probability that our domain is instead the interior of $\Phi_{1}^{-1}(\overline{B_{P}(r_{1})}\cap\Phi_{1}(U_{1}))$ with absorbing boundary $\mathscr{C}_{1}$ and the same target. We can thus localize our domain into interior of $\Phi_{1}^{-1}(\overline{B_{P}(r_{1})}\cap\Phi_{1}(U_{1}))$, which will be denoted by $U^{\prime}$, as shown in figure 5.
@@ 6.2. Explicit Computation and Estimation in $U^{\prime}$
We may now turn to some explicit computations and we restrict the RBM in $U^{\prime}$ for now. Now notice that $\Phi_{1}$ maps $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$ into two semi-spheres by their definition. We denote them by $\mathscr{S}_{1}$, $\mathscr{S}_{2}$, respectively. We also define our RBM in $\Omega$ by $\Phi_{1}$ so we may consider the original RBM satisfies $(1)$ and associated assumptions in $\Phi_{1}(U^{1})$.
@@
Now by adapting the proof in the case of $\mathbb{C}$ with mean value inequality we may assume that there exists constants $k_{1}$, $k_{2}$ depends on $\Phi_{1}$ (hence $\Omega$) such that $B_{P}(k_{1}\epsilon)\cap\mathbb{R}_{0}^{n-1}\subset\Phi_{1}(\mathrm{tg})\subset B_{P}(k_{2}\epsilon)\cap\mathbb{R}_{0}^{n-1}$ for all $\epsilon$ small enough, which allows us to have
@@
(we have abused notation $\tau$). So now we may turn the problem into the asymptotic hitting probability for the following problem: Given the RBM starts at a point $y_{0}$ on $\mathscr{C}_{1}$, with absorbing boundary $\mathscr{C}_{2}$, target at $B_{P}(k_{i}\epsilon)\cap\mathbb{R}_{0}^{n-1}$ for $i=1,2$, and reflecting boundary at $(B_{P}(r_{1})\setminus B_{P}(k_{i}\epsilon))\cap\mathbb{R}_{0}^{n-1}$. The target hitting probability is $\mathbb{P}_{k_{i}}$ for each $i$.
By reflection $\mathbb{P}_{k_{i}}$ is the probability that the Brownian motion $W^{\Phi_{1}\left(B^{x_{0}}_{\tau_{1}}\right)}_{t}$, where $W_{t}$ is stated in equation $(1)$ and restricted in $\overline{B_{P}(r_{1})}\setminus(B_{P}(k_{i}\epsilon)\cap\mathbb{R}_{0}^{n-1})$ with target $B_{P}(k_{i}\epsilon)\cap\mathbb{R}_{0}^{n-1}$ and absorbing boundary $\partial B_{P}(r_{1})$, hits the target at escaping time of $B_{P}(r_{1})$.
@@
$\tau_{2}=\inf\left\{t\geq 0:W^{\Phi_{1}\left(B^{x_{0}}_{\tau_{1}}\right)}_{t}\notin B_{P}(r_{1})\setminus\overline{B_{P}(k_{1}\epsilon)}\right\}$
@@
(5) $\mathbb{P}_{k_{1}}=\mathbb{P}\left(W^{y_{0}}_{\tau_{2}}\in\partial B_{P}(k_{1}\epsilon)\right)\times\mathbb{P}\left(W_{\tau}^{W^{y_{0}}_{\tau_{2}}}\in B_{P}(k_{1}\epsilon)\cap\mathbb{R}_{0}^{n-1}\right).$
Now we observe again that $\mathbb{P}\left(W^{y_{0}}_{\tau_{2}}\in\partial B_{P}(k_{1}\epsilon)\right)$ is the probability of the Brownian Motion $W^{\Phi_{1}\left(B^{x_{0}}_{\tau_{1}}\right)}_{t}$ inside the $n$ dimensional open annulus $B_{P}(r_{1})\setminus\overline{B_{P}(k_{1}\epsilon)}$ with absorbing boundary at $\partial B_{P}(r_{1})$ and target at $\partial B_{P}(k_{1}\epsilon)$. Now we may explicitly write out the solution to this target hitting probability. Since we know that the target hitting probability $\mathbb{P}\left(W^{y_{0}}_{\tau_{2}}\in\partial B_{P}(k_{1}\epsilon)\right)=u_{1}(y_{0})$ satisfies the Dirichlet problem
@@
Now we need to estimate $\mathbb{P}\left(W_{\tau}^{y_{1}}\in B_{P}(k_{1}\epsilon)\cap\mathbb{R}_{0}^{n-1}\right)$ where $W^{y_{0}}_{\tau_{2}}=y_{1}$. We claim the following estimation:
@@
By shifting and rescaling invariance of Brownian motion, let
$Z_{t}:=\frac{H^{y_{1}}_{t}-y_{1}}{4k_{1}\epsilon}$
be another (non standard) Brownian motion started at 00, roams inside $B_{0}(1)$, with original target of $H^{y_{1}}_{t}$ being shifted and rescaled to a $n-1$ dimensional open ball with diameter $\frac{1}{2}$ (which will be denoted by $\mathrm{tg}^{\prime}$) and absorbing boundary being shifted and rescaled to $\partial B_{0}(1)$. Let $\tau_{3}=\inf\{t\geq 0:Z_{t}\notin B_{0}(1)\setminus\mathrm{tg}^{\prime}\}$ Then we have
@@
We notice that $\mathrm{dist}(0,\mathrm{tg}^{\prime})<\displaystyle\frac{1}{4}$. Consider the $n-1$ dimensional cube $Q_{1}$ inscribed in $\mathrm{tg}^{\prime}$ and we notice that there exists a positive constant $c_{4}$ such that it only depends on $n-1$ and the side length of $Q_{1}$ is ${c_{4}}$. Now consider building up another $n$ dimensional cube $Q_{2}$ with one face exactly $Q_{1}$ and 00 lies in same side of $Q_{2}$ compare to $\mathbb{R}^{n-1}_{0}$. Then $Q_{2}$ is well defined and lies inside $B_{0}(1)$. Let the center of $Q_{2}$ be $Q$ and construct sphere $\partial B_{Q}\left(\displaystyle\frac{c_{4}}{4}\right)$, as shown in figure 7.
@@
Let $\mathbb{P}_{3}$ be the probability that $Z_{t}$ hits $\partial B_{Q}\left(\displaystyle\frac{c_{4}}{4}\right)$ before hitting $\partial B_{0}(1)$. Then as $c_{4}$ is independent of $\epsilon$ there exists constant $c_{5}>0$ only depends on $n$ and $k_{1}$ such that $\mathbb{P}_{3}\geq c_{5}$. Let $\mathbb{P}_{4}$ be the probability that starts on a point on $\partial B_{Q}\left(\displaystyle\frac{c_{4}}{4}\right)$ and exits $Q_{2}$ at $Q_{1}$. Then by rescaling invariance we also see there exists $c_{6}>0$ that also only depends on $n$ and $k_{1}$ such that $\mathbb{P}_{4}\geq c_{6}$. So
@@
We may obtain the same result for the case of $k_{2}$.
@@ 6.3. Proof of Main Result
$u_{i}(y_{0})c^{(i)}_{3}\leq\mathbb{P}_{k_{i}}\leq u_{i}(y_{0})$
for $i=1,2$. So by $(2),(3)$ and proposition we proofed in 6.1 we have
@@
We note that $\mathbb{P}\left(\tau_{1}<\infty\right)>0$ as position of $\mathscr{C}_{2}$ is independent of $\epsilon$. We also notice that
So that let $\epsilon\to 0$ we have $u_{i}$ is comparable with general Newtonian potential and thus the main result is proven.

### v4

@@ 6. Proof of Main Result in General SULD Domains
As we remarked, we consider the case when $\epsilon$ is very small. In this case, we can localize our domain
@@ 6.1. Localization of Domain
Observe that when the RBM particle is near the target, the hitting probability is asymptotically irrelevant to the shape of the domain.
Without loss of generality let $\mathrm{tg}\in U^{1}$ and let $\mathscr{C}_{1}$, $\mathscr{C}_{2}$ be two fixed $n-1$ dimensional subsets in $U^{1}$ such that $\Phi_{1}(\mathscr{C}_{1})=\partial B_{P}(r_{1})\cap\Phi_{1}(U_{1})$ and $\Phi_{1}(\mathscr{C}_{2})=\partial B_{P}(r_{2})\cap\Phi_{1}(U_{1})$ for some $r_{1}>r_{2}>0$ and that both curves do not intersects $\mathrm{abs}$. As both $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$ are disjoint compact sets their distance is separated, and we may assume that any curve $\gamma:[0,1]\to\Omega$ connecting some points on $\mathrm{abs}$ and $P=\gamma(1)$ that only intersects $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$ once for each, intersects $\mathscr{C}_{1}$ at $t_{1}$ and $\mathscr{C}_{2}$ at $t_{2}$ where $t_{1}<t_{2}$ ($\mathscr{C}_{2}$ lies “closer” to $P$ than $\mathscr{C}_{1}$). We let $\tau_{1}=\inf\{t\geq 0:B_{t}^{x_{0}}\in\mathscr{C}_{2}\}$. Let $\mathbb{P}_{1}:=\mathbb{P}(B^{x_{0}}_{\tau}\in\mathrm{tg},B^{x_{0}}_{[\tau_{1},\tau]}\cap\mathscr{C}_{1}=\emptyset)$. We claim the following observation:
@@
For $x_{0}$ lies in $B_{P}(r_{1})\cap\Omega\setminus B_{P}(r_{2})$, we have
@@
Since our RBM particle roams outside $\Phi_{1}^{-1}(\overline{B_{P}(r_{1})}\cap\Phi_{1}(U_{1}))$ we have a smaller chance to hit $\mathrm{tg}$ so $\mathbb{P}_{1}\geq\mathbb{P}(B^{x_{0}}_{\tau}\in\mathrm{tg},B^{x_{0}}_{[\tau_{1},\tau]}\cap\mathscr{C}_{1}\neq\emptyset)$. Thus $\mathbb{P}\left(B^{x_{0}}_{\tau}\in\mathrm{tg}\right)\leq 2\mathbb{P}_{1}$. We may take $c_{1}=1$ and $c_{2}=2$. ∎
@@
(3) $\mathbb{P}\left(B^{x_{0}}_{\tau}\in\mathrm{tg}\right)=\mathbb{P}(\tau_{1}<\infty)\times\mathbb{P}\left(B_{\tau}^{B^{x_{0}}_{\tau_{1}}}\in\mathrm{tg}\right)$
as any RBM hits target will hits $\mathscr{C}_{1}$ by geometrical observation. Now from above proposition for asymptotic behavior we may consider $\mathbb{P}\left(B_{\tau}^{B^{x_{0}}_{\tau_{1}}}\in\mathrm{tg},B^{x_{0}}_{[\tau_{1},\tau]}\cap\mathscr{C}_{1}=\emptyset\right)$. This probability can be interpreted as the probability that the RBM starts at a random point in $\mathscr{C}_{2}$, roaming in $\Phi_{1}^{-1}(\overline{B_{P}(r_{1})}\cap\Phi_{1}(U_{1}))$ and hits target, which is same as the target hitting probability that our domain is instead the interior of $\Phi_{1}^{-1}(\overline{B_{P}(r_{1})}\cap\Phi_{1}(U_{1}))$ with absorbing boundary $\mathscr{C}_{1}$ and the same target. We can thus localize our domain into interior of $\Phi_{1}^{-1}(\overline{B_{P}(r_{1})}\cap\Phi_{1}(U_{1}))$, which will be denoted by $U^{\prime}$, as shown in figure 5.
@@ 6.2. Explicit Computation and Estimation in $U^{\prime}$
We may now turn to some explicit computations and we restrict the RBM in $U^{\prime}$ for now. Notice that $\Phi_{1}$ maps $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$ into two semi-spheres by their definition. We denote them by $\mathscr{S}_{1}$, $\mathscr{S}_{2}$, respectively. We also define our RBM in $\Omega$ by $\Phi_{1}$ so we may consider the original RBM satisfies $(1)$ and associated assumptions in $\Phi_{1}(U^{1})$.
@@
Now by adapting the proof in the case of $\mathbb{C}$ with mean value inequality, we may assume that there exists constants $k_{1}$, $k_{2}$ depends on $\Phi_{1}$ (hence $\Omega$) such that $B_{P}(k_{1}\epsilon)\cap\mathbb{R}_{0}^{n-1}\subset\Phi_{1}(\mathrm{tg})\subset B_{P}(k_{2}\epsilon)\cap\mathbb{R}_{0}^{n-1}$ for all $\epsilon$ small enough, which allows us to have
@@
(we have abused notation $\tau$). So now we may turn the problem into the asymptotic hitting probability for the following problem: Given the RBM $X^{y_{0}}_{t}$ starts at a random point $y_{0}$ on $\mathscr{C}_{1}$, with absorbing boundary $\mathscr{C}_{2}$, target at $B_{P}(k_{i}\epsilon)\cap\mathbb{R}_{0}^{n-1}:=\mathrm{tg}_{i}$ for $i=1,2$, and reflecting boundary at $(B_{P}(r_{1})\setminus B_{P}(k_{i}\epsilon))\cap\mathbb{R}_{0}^{n-1}$. The target hitting probability is $\mathbb{P}_{k_{i}}$ for each $i$. The random point $y_{0}$ is distributed according to the hitting density $\mathrm{d}{H(\mathscr{C}_{2})}$ of $B^{x_{0}}_{t}$ at $\mathscr{C}_{2}$ condition on $\tau_{1}<\infty$. Above all, we are interested in
$\int_{x\in\mathscr{C}_{2}}\mathbb{P}\left(\Phi_{1}(B)_{\tau}^{x}\in B_{P}(k_{i}\epsilon)\cap\mathbb{R}_{0}^{n-1},\Phi_{1}(B)_{[\tau_{1},\tau]}^{x_{0}}\cap\mathscr{C}_{1}=\emptyset\right)\mathrm{d}H(\mathscr{C}_{2})$
$=$ $\int_{x\in\mathscr{C}_{2}}\mathbb{P}\left(\Phi_{1}(X)_{\tau}^{x}\in\mathrm{tg}_{i}\right)\mathrm{d}H(\mathscr{C}_{2}).$
We write $\mathbb{P}^{x}_{k_{i}}=\mathbb{P}\left(\Phi_{1}(X)_{\tau}^{x}\in\mathrm{tg}_{i}\right)$.
By reflection $\mathbb{P}^{y_{0}}_{k_{i}}$ is the probability that the Brownian motion $W^{y_{0}}_{t}$, where $W_{t}$ is stated in equation $(1)$ and restricted in $\overline{B_{P}(r_{1})}\setminus(B_{P}(k_{i}\epsilon)\cap\mathbb{R}_{0}^{n-1})$ with target $B_{P}(k_{i}\epsilon)\cap\mathbb{R}_{0}^{n-1}$ and absorbing boundary $\partial B_{P}(r_{1})$, hits the target at escaping time of $B_{P}(r_{1})$.
@@
$\tau_{2}=\inf\left\{t\geq 0:W^{y_{0}}_{t}\notin B_{P}(r_{1})\setminus\overline{B_{P}(k_{1}\epsilon)}\right\}$
@@
(5) $\mathbb{P}^{y_{0}}_{k_{1}}=\mathbb{P}\left(W^{y_{0}}_{\tau_{2}}\in\partial B_{P}(k_{1}\epsilon)\right)\times\mathbb{P}\left(W_{\tau}^{W^{y_{0}}_{\tau_{2}}}\in B_{P}(k_{1}\epsilon)\cap\mathbb{R}_{0}^{n-1}\right).$
Now we observe again that $\mathbb{P}\left(W^{y_{0}}_{\tau_{2}}\in\partial B_{P}(k_{1}\epsilon)\right)$ is the probability of the Brownian Motion $W^{y_{0}}_{t}$ inside the $n$ dimensional open annulus $B_{P}(r_{1})\setminus\overline{B_{P}(k_{1}\epsilon)}$ with absorbing boundary at $\partial B_{P}(r_{1})$ and target at $\partial B_{P}(k_{1}\epsilon)$. Now, we may explicitly write out the solution to this target-hitting probability. Since we know that the target hitting probability $\mathbb{P}\left(W^{y_{0}}_{\tau_{2}}\in\partial B_{P}(k_{1}\epsilon)\right)=u_{1}(y_{0})$ satisfies the Dirichlet problem
@@
We now need to estimate $\mathbb{P}\left(W_{\tau}^{y_{1}}\in B_{P}(k_{1}\epsilon)\cap\mathbb{R}_{0}^{n-1}\right)$ for $y_{1}\in\partial B_{P}(k_{1}\epsilon)$. We claim the following estimation:
@@
By shifting and scaling invariance of Brownian motion, let
$Z_{t}:=\frac{H^{y_{1}}_{(4k_{1}\epsilon)^{2}t}-y_{1}}{4k_{1}\epsilon}$
be another standard Brownian motion started at 00, roams inside $B_{0}(1)$, with original target of $H^{y_{1}}_{t}$ being shifted and scaled to a $n-1$ dimensional open ball with diameter $\frac{1}{2}$ (which will be denoted by $\mathrm{tg}^{\prime}$) and absorbing boundary being shifted and scaled to $\partial B_{0}(1)$. Let $\tau_{3}=\inf\{t\geq 0:Z_{t}\notin B_{0}(1)\setminus\mathrm{tg}^{\prime}\}$ Then we have
@@
We notice that $\mathrm{dist}(0,\mathrm{tg}^{\prime})<\displaystyle\frac{1}{4}$. Consider the $n-1$ dimensional cube $Q_{1}$ inscribed in $\mathrm{tg}^{\prime}$ and we notice that there exists a positive constant $c_{4}$ such that it only depends on $n-1$ and the side length of $Q_{1}$ is ${c_{4}}$. Now consider building up another $n$ dimensional cube $Q_{2}$ with one face exactly $Q_{1}$ and 00 lies on the same side of $Q_{2}$ compared to $\mathbb{R}^{n-1}_{0}$. Then $Q_{2}$ is well defined and lies inside $B_{0}(1)$. Let the center of $Q_{2}$ be $Q$ and construct sphere $\partial B_{Q}\left(\displaystyle\frac{c_{4}}{4}\right)$, as shown in figure 7.
@@
Let $\mathbb{P}_{3}$ be the probability that $Z_{t}$ hits $\partial B_{Q}\left(\displaystyle\frac{c_{4}}{4}\right)$ before hitting $\partial B_{0}(1)$. Then as $c_{4}$ is independent of $\epsilon$ there exists constant $c_{5}>0$ only depends on $n$ and $k_{1}$ such that $\mathbb{P}_{3}\geq c_{5}$. Let $\mathbb{P}_{4}$ be the probability that starts on a point on $\partial B_{Q}\left(\displaystyle\frac{c_{4}}{4}\right)$ and exits $Q_{2}$ at $Q_{1}$. Then by rescaling invariance, we also see there exists $c_{6}>0$ that also only depends on $n$ and $k_{1}$ such that $\mathbb{P}_{4}\geq c_{6}$. So
@@
We may obtain the same result for the case of $k_{2}$. Hence, let the hitting density of $W^{y_{0}}_{t}$ to $\partial B_{P}(k_{1}\epsilon)$ condition on $\tau_{2}<\infty$ to be $\mathrm{d}H_{\epsilon}$. Integrate we have
$\mathbb{P}\left(W_{\tau}^{W^{y_{0}}_{\tau_{2}}}\in B_{P}(k_{i}\epsilon)\cap\mathbb{R}_{0}^{n-1}\right)=$ $\int_{y_{1}\in\partial B_{P}(k_{i}\epsilon)}\mathbb{P}\left(W_{\tau}^{y_{1}}\in B_{P}(k_{i}\epsilon)\cap\mathbb{R}_{0}^{n-1}\right)\mathrm{d}H_{\epsilon}$
$\geq$ $c_{3}^{(i)}\int_{y_{1}\in\partial B_{P}(k_{i}\epsilon)}\mathrm{d}H_{\epsilon}=c_{3}^{(i)}.$
@@ 6.3. Proof of Main Result
$u_{i}(y_{0})c^{(i)}_{3}\leq\mathbb{P}^{y_{0}}_{k_{i}}\leq u_{i}(y_{0})$
for $i=1,2$. We also notice that
So let $\epsilon\to 0$ we have $u_{i}$ is comparable with general Newtonian potential, and $y_{0}$ does not influence the asymptotic behavior. Hence
$\mathbb{P}_{k_{i}}=\int_{x\in\mathscr{C}_{2}}\mathbb{P}\left(\Phi_{1}(X)_{\tau}^{x}\in\mathrm{tg}_{i}\right)\mathrm{d}H(\mathscr{C}_{2})$
is comparable with $u_{i}$. So by $(2),(3)$ and Proposition 6.1 we have
@@
We note that $\mathbb{P}\left(\tau_{1}<\infty\right)>0$ as position of $\mathscr{C}_{2}$ is independent of $\epsilon$. Now, for any general starting point $x\in\Omega$, the probability that it hits $U^{1}$ is a positive constant determined by $x$ and $\Omega$, and hence we arrive at the conclusion.